# Affordance-based Robot Manipulation with Flow Matching

### v4

@@ Abstract
We present a framework for assistive robot manipulation, which focuses on two fundamental challenges: first, efficiently adapting large-scale models to downstream scene affordance understanding tasks, especially in daily living scenarios where gathering multi-task data involving humans requires strenuous effort; second, effectively learning robot action trajectories by grounding the visual affordance model. We tackle the first challenge by employing a parameter-efficient prompt tuning method that prepends learnable text prompts to the frozen vision model to predict manipulation affordances in multi-task scenarios. Then we propose to learn robot action trajectories guided by affordances in a supervised flow matching method. Flow matching represents a robot visuomotor policy as a conditional process of flowing random waypoints to desired robot action trajectories. Finally, we introduce a real-world dataset with 10 tasks across Activities of Daily Living to test our framework. Our extensive evaluation highlights that the proposed prompt tuning method for learning manipulation affordance achieves competitive performance and even outperforms some other finetuning protocols across data scales, while satisfying parameter efficiency. Learning multi-task robot action trajectories with flow matching leads to consistently favorable results in several robot manipulation benchmarks than some alternative behavior cloning methods. This includes more stable training and evaluation, and noticeably faster inference, while maintaining comparable generalization performance to diffusion policy, where flow matching performs marginally better in most cases. Our framework seamlessly unifies affordance learning and action generation with flow matching for robot manipulation. https://hri-eu.github.io/flow-matching-policy/
@@ I Introduction
Recent advances in vision-language models (VLMs) present unprecedented opportunities to solve robot manipulation problems. Attempts in the field have focused on three primary aspects: 1) End-to-end learning manipulation from scratch. These approaches [35] make the least assumptions on tasks and are formulated in language-image-to-action prediction models. 2) Off-the-shelf-vision-language models for robot manipulation. Such line of works have explored using pre-trained VLMs with prompt engineering in various contexts of robot motion learning, including reward design for reinforcement learning [31], python coding [26], joint actions [47], etc. 3) Intermediate substrate to bridge high-level language-image instructions and low-level robot policies. These works usually introduce some form of prior derived from human knowledge as an intermediate stage, including affordances [22], pre-trained visual representations [48], primitive skills [23], etc. In this paper, we follow the third line of work to unify an affordance model and low-level robot policies, which helps to alleviate the sample inefficiency problem of end-to-end learning.
@@
Figure 1: The proposed framework of unifying affordance map learning and action generation for robot manipulation. Given the same visual scene with different language instructions, the model first extracts instruction-relevant manipulation affordances. This is achieved through a prompt tuning method that prepends learnable text-conditioned prompts in a frozen vision foundation model. Then a flow matching policy is proposed to transform the random waypoints to the desired action trajectories, guided by task-relevant affordance maps.
Extracting affordance knowledge has long inspired the robot community [48]. Humans heavily rely on affordances to efficiently perform day-to-day tasks across environments. The concept of affordance has been introduced in [12], referring to the ability to perform certain actions with objects in the context of a given scene. In this work, we first focus on learning affordance maps related to the locations of object areas that permit or enable specific actions. Then, the knowledge of affordance maps is used to plan and execute appropriate physical interactions with objects with our robot learning policy.
@@
We particularly concentrate on multi-task scenarios with text prompting. As shown in Fig. 1, given the same visual scene but with different language instructions, we aim to extract different affordances for robot policy learning through our proposed model. To leverage the ability of pre-trained foundation models while avoiding expensive computational constraints, recent works have explored parameter-efficiently fine-tuning (PEFT) large vision-language models for visual recognition applications [24]. PEFT could be mainly categorized into two main groups: adapter-based methods (e. g., LoRA [20]) and prompt-based methods [29]. One representative line of such research has concentrated on prompt tuning methods, which prepend learnable prompts to the input of a large frozen pre-trained model and optimize them via gradients during finetuning. Studies on randomly-generalized trainable prompts [24] for universal use or condition-admitted prompt variables [43] for better specific task performance have been both explored. It has shown that prompt tuning could match the performance of full finetuning but with substantially less parameter storage in various domains, including visual tracking [53] and cross-domain tasks such as language-dance assessment [52]. Inspired in part by the notion of human cognitive penetrability mechanism [32] that uses linguistic knowledge to tune ongoing visual processing, we target to incorporate learnable text-conditioned prompts into any vision foundation model while keeping it frozen, preserving its visual understanding capabilities, to learn instruction-relevant manipulation affordance maps.
@@
The subsequent challenge involves deploying the visual affordance across robot manipulation learning paradigms. From the traditional behavior cloning with convolutional networks [49] to transformer-based learning structures [42], extensive research has modeled robot action trajectories from visual scenes. A recent line of work builds on successes in diffusion models [3] with Denoising Diffusion Probabilistic Models (DDPM [19]) and Denoising Diffusion Implicit Models (DDIM [44]) to generate motion trajectories to capture multimodal action distributions. Flow Matching is another novel generative method. Sharing theoretical similarities with diffusion process, flow matching aims to regress onto a deterministic vector field to flow samples toward the target distribution. [27] has proven that the simplicity of flow matching objectives allows favorable performance in stable training and generation quality compared to solving complex stochastic differential equations in stochastic denoising diffusion process. We extend flow matching to the robotics domain. As shown in Fig. 1, the proposed method would flow the random waypoints to the desired action trajectories based on multi-task affordances in a single flow matching policy.
@@
We also construct a Real-world Activities of Daily Living (ADLs) dataset with 10 tasks. The novelty of our dataset is that it contains the same scenarios but with multi-task affordance and robot trajectories. Experimental evaluation on our dataset empirically demonstrates that the prompt tuning method for learning affordances achieves performance competitive, and sometimes beyond other finetuning protocols across data scales, vision-language fusion architectures, and prompt variants. Furthermore, we showcase that across several benchmarks, flow matching attains favorable performance in training stability, generation quality, and computational efficiency amongst competing methods of behavior cloning.
This is the first attempt to ground VLM-based affordance with flow matching for real-world robot manipulation. We have also first systematically evaluated robot manipulation with flow matching on several benchmarks, including various input representations, robot control types, and manipulation tasks. Our code has been released https://github.com/HRI-EU/flow_matching.
@@
    A parameter-efficient prompt tuning method for adapting pretrained vision foundation model conditioned on language instructions to learn manipulation affordances.
@@
    A novel formulation using flow matching for closed-loop 6D robot manipulation learning from various inputs, including visual affordances.
@@
    Empirical and extensive results show that flow matching leads to consistently favorable results than some alternative behavior cloning methods in various manipulation tasks. This includes more stable training and evaluation, and noticeably faster inference than diffusion policy, while maintaining comparable generalization performance to DDIM and slightly better performance than DDPM.
Note that our goal is not to achieve the state-of-the-art general robot manipulation performance, but instead to broadly explore a new paradigm of efficiently adapting VLMs for affordance learning, and robot policy for multimodal action distributions.
@@ II Related Work
II-A Robot Learning from Demonstration
Imitation learning has been a common paradigm for robots which requires simulated or real-world demonstration data collection [50]. To improve data efficiency, extensive works have been proposed to learn robot policies on the top of visual representations [28], such as keypoints or affordance heatmaps [28], instead of end-to-end raw images [13]. This paper concentrates on using affordances to guide the low-level robot manipulation. In terms of network architectures for robot learning, prior works have successfully investigated convolutional networks [49], Transformers [45], generative adversarial networks [18], Energy-Based Models [7], etc. However, the collected data is usually expected to be non-convex and multi-modal due to the variability in human demonstrations. Recent works have addressed this problem by reformulating the robot policy as a generative process. Diffusion policy [3] has emerged as a powerful class of generative models for behavior cloning by representing a robot’s visuomotor policy as a conditional denoising diffusion process. In this work, we investigate flow matching [27], a novel generative model that has demonstrated its superiority in image generation, but is much less explored in robotics domains.
@@ II-B Parameter-Efficient Finetuning
Instruction-aware vision encoding [15] has been extensively studied for language-vision fusing tasks [39]. Given the dominance of large-scale vision-language models, many approaches have been proposed to efficiently finetune a frozen pretrained model for downstream tasks to speed up training and reduce memory. Two representative methods among them are adapters and prompting. The first line of research varies depending on the adapter that could add extra lightweight modules [10, 20]. Other work focuses on prompting [29], which originally primes a frozen pretrained language model for downstream tasks by including a hard text prompt. Recent works on prompt tuning [30] treat soft prompts as continuous vectors and compute their gradients with backpropagation during training. The extension of prompt tuning to vision tasks has gained massive success. Visual prompt tuning [24] has manipulated visual prompts to steer models in arbitrary vision tasks.  [43] have explored extending the random learnable prompts to condition-based prompts that are less universal but more accurate. Inspired by its recent success, we extend the prompt tuning technologies to address the challenge of adapting large pretrained vision-language models to affordance learning for robot manipulation. The intuition is clear: if the model understands the posed text instruction and the inherent context, it should extract visual affordances that directly correspond to the relevant image aspects. Our method achieves the above goal by integrating learnable text-conditioned prompts into a large vision encoder, while keeping it frozen to preserve visual understanding capabilities.
II-C Flow Matching in Robotics
@@
Despite its recent progress in image generation [1], the application of flow matching in robotics domains remains underexplored [21, 41]. Few prior studies have concentrated exclusively on certain robot scenarios for deploying flow matching, for pointcloud environment [4], Riemannian manifold [5], etc. We propose to use flow matching to learn multi-task robot behaviors from raw observations, including visual affordances obtained from a vision-language model, in a single supervised policy. We have first systematically evaluated robot manipulation with flow matching on several benchmarks, including various input representations, robot control types, and manipulation tasks.
@@ III-A Prompt Tuning for Affordance Map Learning
Providing any type of pre-trained vision transformer, our objective is to learn a set of text-conditioned prompts to maximize the likelihood of correct affordance labels, as shown in Fig. 2. Only the prompt-related layers and the decoder are being updated during the training, while the vision transformer remains frozen. Inspired by Vision Prompt Tuning [24], we propose shallow and deep network architectures.
Figure 2: Overview of prompt tuning structures used for affordance learning. (Left) For the shallow structure, text-conditioned prompts are prepended to the first vision transformer layer. (Right) For the deep structure, prompts are inserted to every vision layer. Only the prompt-related layers and the decoder are being updated during the training, while the vision transformer remains frozen.
@@ III-A2 Deep Architecture
[_$[\_,\bm{E}_{1}]$ $=\bm{L}_{1}^{v}([\bm{P}_{1},\bm{E}_{0}])$
[_$[\_,\bm{E}_{i}]$ $=\bm{L}_{i}^{v}([\bm{P}_{i},\bm{E}_{i-1}])$
Figure 3: Framework of flow matching policy. (a) General formulation. At each time step, flow matching takes visual observation $\bm{o}$ (e. g., state-based inputs, RGB-D images, visual affordances) as input, and outputs robot actions (e. g., 6D robot end-effector trajectories, robot joint actions, gripper actions). (b-d) Visualization of the inference process of transforming random waypoints to target actions over time from 0 (green) to 1 (purple). Green lines in (b) denote the flow paths. Vector fields are shown in (c).
@@ III-A3 Implementation Details
Our goal is to integrate textual representations into any vision encoder while keeping it frozen, preserving its visual understanding capabilities. Thus, we have chosen the most basic vision backbone, a pretrained ViT-B-16 transformer. The text layers adopt the classic CLIP setup [39], including Multiheaded Self-Attention, Feed-Forward Networks with LayerNorm and residual connections, and output 76 tokens. As suggested by MAE [17], the decoder is only used for downstream tasks and could be flexible and lightweight. Thus we use one single transformer decoder layer.
We adopt the L2 Mean Squared loss between the predicted and ground truth affordances for network training. The training parameters for the prompt tuning network include an image size of 224 × 224, AdamW with a learning rate of 1.5e-4 including Warmup with step-decay, and batch size of 256. We add positional embeddings to all the image and language tokens to preserve the positional information. In the subsequent experiments, we will further ablate multiple model variants, including text and vision fusion structures, prompt depth, pretrained weights for vision transformers, etc.

### v5

@@ Abstract
We present a framework for assistive robot manipulation that addresses two fundamental challenges: efficient adaptation of large-scale models for scene affordance understanding and effective learning of robot actions by grounding the visual affordance. To tackle the first challenge, we adopt a parameter-efficient prompt tuning method, prepending learnable text prompts to a frozen vision model to predict affordances, while considering spatial and semantic relationships in multi-task scenarios. For the second challenge, we propose a flow matching method, representing a robot visuomotor policy as a conditional process of flowing random waypoints to desired robot actions. We introduce a real-world dataset with 10 tasks to evaluate our approach. Experiments show our prompt tuning method achieves competitive or superior performance to other finetuning protocols across data scales, while satisfying parameter efficiency. Flow matching yields more stable training and faster inference, while maintaining comparable generalization performance to diffusion policy. Our framework seamlessly unifies parameter-efficient affordance learning and robot action generation with flow matching. https://hri-eu.github.io/flow-matching-policy/
@@ I Introduction
Recent advances in vision-language models (VLMs) present unprecedented opportunities to solve robot manipulation problems. Attempts in the field have focused on three primary aspects: 1) End-to-end learning manipulation from scratch. These approaches [padalkar2023open] make the fewest assumptions on tasks and are formulated in language-image-to-action prediction models. 2) Off-the-shelf-vision-language models for robot manipulation. These works have prompted pre-trained VLMs in various contexts of robot motion learning, including reward design for reinforcement learning [ma2023eureka], python coding [liang2023code], joint actions [wang2023prompt], etc. 3) Intermediate substrate to bridge high-level language-image instructions and low-level robot policies. These works usually introduce some form of prior derived from human knowledge as an intermediate stage to alleviate the sample inefficiency problem of end-to-end learning, including affordances [huang2023voxposer], primitive skills [ingelhag2024robotic], etc. In this paper, we follow the third line of work to unify a parameter-efficient affordance model and a low-level robot flow matching policy.
@@
Figure 1: The proposed framework of unifying affordance map learning and action generation for robot manipulation. Given the same visual scene with different language instructions, the model first extracts instruction-relevant manipulation affordances. This is achieved through a prompt tuning method that prepends learnable text-conditioned prompts in a frozen vision foundation model. Then, a flow matching policy is proposed to transform the random waypoints to the desired action trajectories, guided by task-relevant affordance maps.
Extracting affordance knowledge has long inspired the robot community [xiao2022masked]. Recent state-of-the-art works have proposed the affordance-based robot policy by describing affordances with 6D poses or 2D point tracks [nasiriany2024rt, bharadhwaj2024track2act]. For example,  [nasiriany2024rt] represents affordances as the relative poses between the robot end effector and the object at key stages of the task.  [bharadhwaj2024track2act] uses 2D keypoint track predictions to infer a sequence of rigid transforms of the object affordance to be manipulated. Other research decouples affordance learning and robot policy by using more straightforward affordance representations like 2D masks or keypoints [stone2023open, yuan2024robopoint]. These works typically localize task-specific affordances directly on objects relevant to certain actions, given detailed text instructions. Our affordance model additionally considers the relational context needed for interactions, including the spatial and semantic functional interplay relationship between multiple elements. We concentrate on multi-task scenarios with text prompting. As shown in Fig. 1, given the same visual scene but with different language instructions, we aim to extract different affordances for robot policy learning through our proposed model. For example, given a simple instruction of ‘feeding the human’, a sequence of 2D affordance heatmaps on a spoon handle, food and a person’s mouth are identified.
@@
Recent affordance models have successfully fine-tuned pre-trained vision-language foundation models to extract affordance knowledge [qian2024affordancellm, yuan2024robopoint]. To leverage the capabilities of pre-trained foundation models while simultaneously mitigating the associated computational costs, some works have explored parameter-efficient fine-tuning (PEFT) large vision-language models [jia2022visual]. One representative line of PEFT research has concentrated on prompt tuning methods, which prepend learnable prompts to a large frozen pre-trained model and optimize them via gradients during finetuning. In contrast to studies in NLP, it has been shown that prompt tuning could match or even outperform full fine-tuning and adapter-based methods, but with substantially less parameter storage in visual domains [zhu2023visual]. Inspired in part by the notion of human cognitive penetrability mechanism [maier2019no] that uses linguistic knowledge to tune ongoing visual processing, we aim to incorporate learnable text-conditioned prompts into any vision foundation model while keeping it frozen, preserving its visual understanding capabilities, to learn instruction-relevant manipulation affordance maps.
@@
The subsequent challenge involves deploying the visual affordance across robot manipulation learning paradigms. A recent line of work builds on successes in curriculum-based imitation learning [blessing2023information], probabilistic movement primitive modeling [yildirim2024conditional], and diffusion models [chi2023diffusion] to generate motion trajectories to capture multimodal action distributions. Flow Matching is another novel generative method. Sharing theoretical similarities with diffusion process, flow matching aims to regress onto a deterministic vector field to flow samples toward the target distribution. [lipman2022flow] has proven that the simplicity of flow matching objectives allows favorable performance in stable training and generation quality compared to solving complex stochastic differential equations in the stochastic denoising diffusion process. We extend flow matching to the robotics domain. As shown in Fig. 1, the proposed method would flow the random waypoints to the desired action trajectories based on multi-task affordances in a single flow matching policy.
@@
We also construct a Real-world Activities of Daily Living (ADLs) dataset with 10 tasks. The novelty of our dataset is that it contains the same scenarios but with multi-task affordance and robot trajectories. Experimental evaluation on our dataset empirically demonstrates that the prompt tuning method for learning affordances achieves performance competitive, and sometimes beyond other finetuning protocols across data scales and vision-language fusion architectures.
This work seamlessly grounds VLM-based affordance with flow matching for real-world robot manipulation. We leverage the capability of the flow matching policy to represent multi-modal action distributions for learning accurate 6D robot actions, while 2D affordance maps readily provide sufficient guidance in shaping the policy. Depending on the system design, our approach can run at approximately 30 Hz (affordance learning: 18.340 ms; flow matching with one-step inference: 13.228 ms), making it suitable for closed-loop manipulation tasks. We have further systematically evaluated robot manipulation with flow matching on several benchmarks, including various input representations, robot control types, and manipulation tasks. We showcase that across several benchmarks, flow matching attains favorable performance in training stability, generation quality, and computational efficiency amongst competing methods of behavior cloning. Our code is publicly available.
@@
    A parameter-efficient prompt tuning method for adapting pretrained vision foundation model conditioned on language instructions to learn manipulation affordances, incorporating the spatial and semantic interactions between elements in multi-task scenarios.
@@
    A novel formulation using flow matching for closed-loop 6D robot manipulation learning from various inputs, including visual affordances. Empirical and extensive results show that flow matching leads to consistently favorable results than some alternative behavior cloning methods in various manipulation tasks. This includes more stable training and evaluation and noticeably faster inference, while maintaining comparable generalization performance to diffusion policy (DDIM, DDPM and score-based models).
@@
    A framework seamlessly unifies parameter-efficient affordance learning and robot action generation with flow matching. Experimental results, compared against end-to-end learning and off-the-shelf VLM-based trajectory generation frameworks, demonstrate that affordance representations provide consistent guidance for shaping manipulation policies.
@@ II Related Work
II-A Affordance Learning and Affordance-Based Robot Policy
Humans rely heavily on affordances to perform day-to-day tasks across environments efficiently. The concept of affordance has been introduced in [gibson2014ecological], referring to the ability to perform certain actions with objects in the context of a given scene. Several state-of-the-art works have adopted this concept and successfully proposed affordance-based robot policy by learning the affordance parameterization that could be amenable to deployment on robots, including 6D pose [nasiriany2024rt], stable object placement [zeng2021transporter], point tracks [bharadhwaj2024track2act], and post-contact trajectory [bahl2023affordances]. Other approaches decouple actionable representation for affordances from robot policy by using more straightforward affordance representations, like 2D masks [stone2023open], keypoints [manuelli2019kpam], heatmaps [zhang2022learning], part segmentation [do2018affordancenet], dense image feature descriptors [florence2018dense]. We follow this idea of decoupling and use 2D heatmaps as affordances.
Recent SoA works have successfully leveraged pre-trained vision-language models (e. g., LLaVA and Vicuna) to extract affordance knowledge, as in AffordanceLLM [qian2024affordancellm] and Robopoint [yuan2024robopoint]. This raises three important research questions: i) How can language-vision inputs for instruction-aware vision encoding be readily fused? ii) How to parameter-efficiently finetune a pretrained foundation model? and iii) How to encode the spatial and semantic functional interplay relationship between elements in multi-task scenarios for robot manipulation affordance learning? We introduce a prompt tuning method as a solution to these questions.
@@ II-B Parameter-Efficient Finetuning
Given the dominance of large-scale vision-language models, many approaches have been proposed to efficiently finetune a frozen pretrained model for downstream tasks to speed up training and reduce memory [gupta2022towards, radford2021learning]. Two representative parameter-efficient finetuning methods are adapters and prompting. The adapter-based research designs the adapter variants that could add extra lightweight modules [gao2024clip, hu2021lora, liu2023tail, sharma2023lossless]. Other work focuses on prompt tuning [liu2021p, liu2023pre], which treats learnable prompts as continuous vectors and computes their gradients with backpropagation during training. Studies on randomly generalised trainable prompts [jia2022visual] for universal use or condition-admitted prompt variables [sohn2023visual] for better specific task performance have both been explored. The extension of prompt tuning to vision tasks has gained massive success. Visual prompt tuning [jia2022visual] has manipulated visual prompts to steer models in arbitrary vision tasks. Dancemvp [zhong2024dancemvp, zhong2023contrastive] has adjusted language prompts to guide text-audio models in performing dancing assessment tasks. Inspired by its recent success, we extend the prompt tuning technologies to address the challenge of adapting large pretrained vision-language models to affordance learning for robot manipulation. The intuition is clear: if the model understands the posed text instruction and the inherent context, it should extract visual affordances that directly correspond to the relevant image aspects. Our method achieves the above goal by integrating learnable text-conditioned prompts into a large vision encoder, while keeping it frozen to preserve visual understanding capabilities. Besides, recent research on VLM for affordance-based robot policy typically localizes task-specific affordances directly on objects relevant to certain actions, which requires detailed language instructions about all task-related objects and actions. We aim to enable our affordance model to incorporate relational context essential for interactions, capturing both spatial and semantic-functional relationships among multiple objects and persons, where only simple prompts are required.
II-C Robot Learning from Demonstration
Imitation learning has been a common paradigm for robots, which requires simulated or real-world demonstration data collection [levine2016end]. To improve data efficiency, extensive work has been proposed to learn robot policies on top of visual representations [liu2024moka], such as keypoints or affordance heatmaps [liu2024moka], instead of end-to-end raw images [goyal2023rvt]. This paper concentrates on using affordances to guide low-level robot manipulation. In terms of network architectures for robot learning, prior works have successfully investigated convolutional networks [levine2016end], Transformers [sridhar2023memory], generative adversarial networks [ho2016generative], Energy-Based Models [florence2022implicit], etc. However, the collected data is usually expected to be non-convex and multi-modal due to the variability in human demonstrations. Recent works have addressed this problem by reformulating the robot policy as a generative process. Diffusion policy [chi2023diffusion], including Denoising Diffusion Probabilistic Models (DDPM [ho2020denoising]), Denoising Diffusion Implicit Models (DDIM [song2020denoising]) and score-based models [reuss2023goal], has emerged as a powerful class of generative models for behavior cloning by representing a robot’s visuomotor policy as a conditional denoising diffusion process. In this work, we investigate flow matching [lipman2022flow], a novel generative model that has demonstrated its superiority in image generation, but is much less explored in robotics domains.
II-D Flow Matching in Robotics
@@
Despite its recent progress in image generation [albergo2022building], the application of flow matching in robotics domains remains underexplored [hu2024adaflow, rouxel2024flow]. Few prior studies have concentrated exclusively on certain robot scenarios for deploying flow matching, for pointcloud environment [chisari2024learning], dynamics [nguyen2025flowmp], Riemannian manifolds [braun2024riemannian], etc. We propose to use flow matching to learn multi-task robot behaviors from raw observations, including visual affordances obtained from a vision-language model, in a single supervised policy. We have first systematically evaluated robot manipulation with flow matching on several benchmarks, including various input representations, robot control types, and manipulation tasks.
@@ III-A Prompt Tuning for Affordance Map Learning
Providing any type of pre-trained vision transformer, our objective is to learn a set of text-conditioned prompts to maximize the likelihood of correct affordance labels, as shown in Fig. 2. Only the prompt-related layers and the decoder are being updated during the training, while the vision transformer remains frozen. Inspired by Vision Prompt Tuning [jia2022visual], we propose two frameworks: shallow and deep network architectures.
Figure 2: Overview of prompt tuning structures used for affordance learning. (Left) For the shallow structure, text-conditioned prompts are prepended to the first vision transformer layer. (Right) For the deep structure, prompts are inserted into every vision layer. Only the prompt-related layers and the decoder are being updated during the training, while the vision transformer remains frozen.
@@ III-A2 Deep Architecture
$[\_,\bm{E}_{1}]$ $=\bm{L}_{1}^{v}([\bm{P}_{1},\bm{E}_{0}])$
$[\_,\bm{E}_{i}]$ $=\bm{L}_{i}^{v}([\bm{P}_{i},\bm{E}_{i-1}])$
This is different to parallel adapters, which introduce additional trainable modules that run in parallel with the main transformer layers instead of modifying existing layers.
Figure 3: Framework of flow matching policy. (a) General formulation. At each time step, flow matching takes visual observation $\bm{o}$ (e. g., state-based inputs, RGB-D images, visual affordances) as input, and outputs robot actions (e. g., 6D robot end-effector trajectories, robot joint actions, gripper actions). (b-d) Visualization of the inference process of transforming random waypoints to target actions over time from 0 (green) to 1 (purple). Red lines in (b) denote the flow paths. Vector fields are shown in (c).
@@ III-A3 Implementation Details
In our implementation, we select and adopt a single structure from the two designs. The deep structure demonstrates better performance than the shallow structure, albeit with increased computational cost. Our goal is to integrate textual representations into any vision encoder while keeping it frozen, preserving its visual understanding capabilities. Thus, we have chosen the most basic vision backbone, a pretrained ViT-B-16 transformer. We use the classic CLIP Transformer layers [radford2021learning] that output 76 tokens. As suggested by MAE [he2022masked], the decoder is only used for downstream tasks and could be flexible and lightweight. Thus we use one single transformer decoder layer.
We use the L2 Mean Squared loss between the predicted and ground truth affordances for network training. The training parameters for the prompt tuning network include an image size of 224 × 224, AdamW with a learning rate of 1.5e-4, including Warmup with step-decay, and a batch size of 256. We add positional embeddings to all the image and language tokens to preserve the positional information. In the subsequent experiments, we will further ablate multiple model variants, including text and vision fusion structures, prompt depth, pretrained weights for vision transformers, etc.